I do not understand how to find the point slope form of the equation " Through: (-3,-3), and parallel to y=7/3x+3."
Answer by rfer(16322)

-3=7/3*-3+b
-3=-21/3+b
-3=-7+b
-3+7=b
4=b
-----------------
y=7/3x+7